The following labor standards have been established for a particular product: Standard labor-hours per unit of output 9.6 hours Standard labor rate $ 13.40 per hour The following data pertain to operations concerning the product for the last month: Actual hours worked 7,400 hours Actual total labor cost $ 96,200 Actual output 950 units What is the labor efficiency variance for the month?

Multiple Choice

$26,008 F

$26,008 U

$22,360 F

$23,048 F
Answer #1

D. $23,048 F

Standard hours = Standard hours per unit × Actual output = 9.6 × 950 = 9,120

Labor efficiency variance = Standard rate × (Actual hours − Standard hours)

= $13.40 × (7,400 − 9,120) = $23,048 favorable